Operations Report – Mar 6th

Crew 206 Operations Report 06 – MAR – 2019

SOL: 10

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and
alternator need replacement on generator. A contractor has
been arranged to make these repairs within a few days.

Generator (hours run): 13 hours

Solar— SOC % – Turned off Gen : 80%, generator still off at 1900 (87%).

SOC tracking : Tracking is a mess today because of the special events. I
will get a cleaner baseline tomorrow.

Notes on electricity: We have no special needs in electricity other than
some computers and heavy batteries charging.

Diesel Reading – 65%
Propane Reading – 79%

Ethanol Free Gasoline – 7 gallons

Water (Auxillary tank) – Not in use
Water (Static tank) – 410 gallons
Water differential (static tank) – 11 gallons

Gallons transferred: Not Applicable

Water in GreenHab – 241 gallons

Static to Loft Pump used – Yes

Water Meter: 0141677,1 gallons

Toilet tank emptied: Yes

Deimos rover used: No, still not functional
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A
Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Not Applicable
Beginning charge: Not Applicable
Ending charge: Not Applicable
Currently charging: Not Applicable

Opportunity rover used: Not used
Hours: 53,0 hours
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: Used
Hours: 77,8 hours
Beginning charge: 100%
Ending charge: 74%
Currently charging: Yes

Spirit rover used: Used
Hours: 78,9 hours
Beginning charge: 100%
Ending charge: 56%
Currently charging: Yes

Notes on rovers: Nothing to report.

ATV’s Used: N/A
Reason for use: Not used
Oil Added? N/A
ATV Fuel Used: N/A
# Hours the ATVs were Used today: None
Notes on ATVs: None

Hab Car used and why, where? Not used.

Staff Crew Car used and why, where? Off Campus

General notes and comments: Nothing to report.

Summary of internet: Still investigating. Nothing more to report today.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report.

Summary of Science Dome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing more to report.